# Intuitive Fine-Tuning: Towards Simplifying Alignment into a Single Process

### v2

@@ A.3 Proof for Bellman Equation
$\exp\big{(}-\mathcal{L}(\hat{T_{\theta}}(s_{n}^{*},\rho_{0}))\big{)}$ $=\mathcal{T}_{\theta}(a_{n}^{*},\delta_{\theta}(s_{n}^{*}))\bigg{(}\sum\limits_{n+1}^{N}\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))\bigg{)}$ (26)
@@ A.4 Reformulation of Typical Methods
where the human’s preference is unbiasedly estimated, but the model’s preference is inaccurately represented by using $s_{i}^{*}$.
@@
$\mathcal{L_{R}}=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{i}^{+}\sim\mathcal{S}_{\rho_{0}}^{+},s_{i}^{-}\sim\mathcal{S}_{\rho_{0}}^{-}}\left[-\log\sigma\bigg{(}\sum\limits_{i=0}^{N}\log\mathcal{T_{R}}(\pi^{+}(s_{i}^{+})|s_{i}^{+})-\sum\limits_{i=0}^{N}\log\mathcal{T_{R}}(\pi^{-}(s_{i}^{-})|s_{i}^{-})\bigg{)}\right]$ (30)
@@
$\mathcal{L}_{\text{DPO}}=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{i}^{*}\sim\mathcal{S}_{\rho_{0}}^{*},s_{i}^{\theta}\sim\mathcal{S}_{\rho_{0}}^{\theta}}\left[-\log\sigma\bigg{(}\sum\limits_{i=0}^{N}\log\mathcal{T}_{\theta}(\pi^{*}(s_{i}^{*}),s_{i}^{*})-\sum\limits_{i=0}^{N}\log\mathcal{T}_{\theta}(\pi_{\theta}(s_{i}^{\theta}),s_{i}^{\theta})\bigg{)}\right]$ (31)
@@
$\mathcal{L}_{\text{DPO}}=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{i}^{+}\sim\mathcal{S}_{\rho_{0}}^{+},s_{i}^{-}\sim\mathcal{S}_{\rho_{0}}^{-}}\left[-\log\sigma\bigg{(}\sum\limits_{i=0}^{N}\log\mathcal{T}_{\theta}(\pi^{+}(s_{i}^{+}),s_{i}^{+})-\sum\limits_{i=0}^{N}\log\mathcal{T}_{\theta}(\pi^{-}(s_{i}^{-}),s_{i}^{-})\bigg{)}\right]$ (32)
@@
$\mathcal{L}_{\text{IFT}}=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{i}^{*}\sim\mathcal{S}_{\rho_{0}}^{*}}\left[-\sum\limits_{n=0}^{N}\sum\limits_{i=n}^{N}\log\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))\right]$ (33)
@@ B.1 NLP Settings
For the coefficient $\beta$ in DPO and ORPO, we use 0.1 and 0.25 respectively, as presented in their original papers. For IFT, we choose 0.2 for $\lambda$ and incorporate a decay factor of 0.95 to fitting better with the Bellman Equation. We save checkpoints every 20k steps and select the results from the checkpoint with the best average score to demonstrate the performance of each method.
@@
         Name          Value
         epoch          3
         mini batch size          8
         gradient accumulation step          64
         warmup ratio          0.1
         scheduler          cosine
         learning rate          5e-7
         optimizer          RMSprop
         precision          bfloat16
@@ B.2 Frozen Lake Setting
B.3 Pseudo-code of IFT
Algorithm 1 Intuitive Fine-Tuning
1:  Input:     Initial instruction $\rho_{0}$     Ground truth answer $s^{*}$ with $N$ tokens, $s^{*}[1],\ldots,s^{*}[N]$
2:  Step 1: Inference One Step Ahead
3:  for $t$ in $[1,N]$ do
5:     Sample the predicted token: $s^{\theta}[t]=\arg\max P_{t}^{\prime}$
6:  end for
7:  Step 2: Intuitive Preference Estimation
8:  Encode $s^{*}$ and $s^{\theta}$ using the Embedding Layer $E$
9:  Compute the fused embedding $e=(1-\lambda)E(s^{*})+\lambda E(s^{\theta})$
10:  for $t$ in $[1,N]$ do
12:     Compute the token-level loss: $\mathcal{L}_{t}=\log(P_{t}^{\prime\prime},s^{*}[t])$
13:  end for
14:  Step 3: Dynamic Relation Propagation
15:  for $t$ in $[1,N]$ do
17:  end for
18:  Output: Final loss $\mathcal{L}_{\text{IFT}}=w\cdot\mathcal{L}$
1def get_batch_logps(logits,
2 labels):
3
4 labels = labels[:, 1:].clone()
5 logits = logits[:, :-1, :]
6 loss_mask = (labels != -100)
7
8 labels[labels == -100] = 0
9
10 per_token_logps = torch.gather(logits.log_softmax(-1),
11 dim=2,
12 index=labels.unsqueeze(2)
13 ).squeeze(2)
14
15 return per_token_logps * loss_mask
1def get_ift_loss(policy,
2 input_ids,
3 attention_mask,
4 labels,
5 _lambda: float=0.2,
6 propagation_decay: float=0.95):
7
8 from torch_discounted_cumsum import discounted_cumsum_right
9 loss_mask = (labels[:, 1:] != -100)
10
11 # Step 1: Inference One Step Ahead
12 inputs_embeds = policy.embed_tokens(input_ids)
13 with torch.no_grad():
14 logits = policy(inputs_embeds=inputs_embeds,
15 attention_mask=attention_mask).logits
16 logps = get_batch_logps(logits, labels)
17
18 tokens_further = torch.cat((input_ids[:, 0].unsqueeze(-1),
19 torch.argmax(logits, dim=-1)[:, :-1]),
20 dim=-1)
21 input_ids_further = torch.where(labels==-100,
22 input_ids,
23 tokens_further)
24 attention_mask_further = attention_mask
25
26 # Step 2: Intuitive Preference Estimation
27 inputs_embeds_further = self.embed_tokens(input_ids_further)
28 inputs_embeds_further = (1 - _lambda) * inputs_embeds \
29 + _lambda * inputs_embeds_further
30 logits_further = policy(inputs_embeds=inputs_embeds_further,
31 attention_mask=attention_mask_further,
32 ).logits
33 logps_further = _get_batch_logps(logits_further, labels)
34 loss_orig = -logps / loss_mask.sum(-1).unsqueeze(-1)
35
36 # Step 3: Dynamic Relation Propagation
37 cumsum_weight = discounted_cumsum_right(loss_orig,
38 gamma=propagation_decay)
39 cumsum_weight[loss_mask == 0] = 0
40 cumsum_weight *= (loss_mask.sum(-1, keepdim=True) \
41 / cumsum_weight.sum(-1, keepdim=True))
42
43 loss_ift = (-logps * cumsum_weight).sum(-1) / loss_mask.sum(-1)
44
45 return loss_ift.mean()

### v3

@@ A.3 Proof for Bellman Equation
$\exp\big{(}-\mathcal{L}(\hat{T_{\theta}}(s_{n}^{*},\rho_{0}))\big{)}$ (29)
$=\mathcal{T}_{\theta}(a_{n}^{*},\delta_{\theta}(s_{n}^{*}))\bigg{(}\sum\limits_{n+1}^{N}\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))\bigg{)}$
@@ A.4 Reformulation of Typical Methods
where the human’s preference is unbiasedly estimated, but the model’s preference is inaccurately represented by $s_{i}^{*}$.
@@
$\mathcal{L_{R}}$ $=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{i}^{+}\sim\mathcal{S}_{\rho_{0}}^{+},s_{i}^{-}\sim\mathcal{S}_{\rho_{0}}^{-}}$ (33)
$\left[-\log\sigma\bigg{(}\sum\limits_{i=0}^{N}\log\mathcal{T_{R}}(\pi^{+}(s_{i}^{+})|s_{i}^{+})\right.$
$\left.\quad\quad\quad\quad\quad-\sum\limits_{i=0}^{N}\log\mathcal{T_{R}}(\pi^{-}(s_{i}^{-})|s_{i}^{-})\bigg{)}\right]$
@@
$\mathcal{L}_{\text{DPO}}$ $=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{i}^{*}\sim\mathcal{S}_{\rho_{0}}^{*},s_{i}^{\theta}\sim\mathcal{S}_{\rho_{0}}^{\theta}}$ (34)
$\left[-\log\sigma\bigg{(}\sum\limits_{i=0}^{N}\log\mathcal{T}_{\theta}(\pi^{*}(s_{i}^{*}),s_{i}^{*})\right.$
$\left.\quad\quad\quad\quad\quad-\sum\limits_{i=0}^{N}\log\mathcal{T}_{\theta}(\pi_{\theta}(s_{i}^{\theta}),s_{i}^{\theta})\bigg{)}\right]$
@@
$\mathcal{L}_{\text{DPO}}$ $=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{i}^{+}\sim\mathcal{S}_{\rho_{0}}^{+},s_{i}^{-}\sim\mathcal{S}_{\rho_{0}}^{-}}$ (35)
$\left[-\log\sigma\bigg{(}\sum\limits_{i=0}^{N}\log\mathcal{T}_{\theta}(\pi^{+}(s_{i}^{+}),s_{i}^{+})\right.$
$\left.\quad\quad\quad\quad\quad-\sum\limits_{i=0}^{N}\log\mathcal{T}_{\theta}(\pi^{-}(s_{i}^{-}),s_{i}^{-})\bigg{)}\right]$
@@
$\mathcal{L}_{\text{IFT}}$ $=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{i}^{*}\sim\mathcal{S}_{\rho_{0}}^{*}}$ (36)
$\left[-\sum\limits_{n=0}^{N}\sum\limits_{i=n}^{N}\log\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))\right]$
@@ B.1 NLP Settings
For the coefficient $\beta$ in DPO, TDPO, ORPO and SimPO, we use 0.1, 0.1, 0.25 and 2.0 respectively, as presented in their original papers. For the coefficient $\gamma/\beta$ ration in SimPO, we use 0.8 to keep the same setting in its original papers. For IFT, we choose 0.2 for $\lambda$ and incorporate a decay factor of 0.95 to fitting better with the Bellman Equation. We save checkpoints every 20k steps and select the results from the checkpoint with the best average score to demonstrate the performance of each method.
         Name          Value          epoch          3          mini batch size          8          gradient accumulation step          64          warmup ratio          0.1          scheduler          cosine          learning rate          5e-7          optimizer          RMSprop          precision          bfloat16
@@ B.2 Frozen Lake Setting
Appendix C More Experimental Results
Method ARC ARC-Gen MMLU TruthfulQA WinoGrande GSM8K Avg. Gemma-2B 42.75 43.17 35.68 35.25 66.46 16.98 39.42      fine-tuning with Gemma2-UltraFeedback-armnorm-60k + SFT 42.06 42.75 34.30 41.49 64.88 21.53 40.85 + DPO 41.30 40.61 35.47 30.11 65.51 18.95 38.27 + TDPO 41.21 40.70 35.62 31.33 65.04 18.88 38.42 + ORPO 41.89 42.06 36.43 41.98 65.90 20.54 41.35 + SimPO 41.38 40.10 35.32 28.76 65.27 20.39 38.22 + IFT 42.49 42.66 35.77 45.41 66.14 22.14 42.39 LLaMA-3-8B 49.40 73.89 62.17 46.63 76.80 50.26 57.05      fine-tuning with LLaMA3-UltraFeedback-armnorm-60k + SFT 52.83 75.00 63.24 50.42 76.95 51.09 58.91 + DPO 51.19 74.23 62.21 36.35 76.24 51.25 55.45 + TDPO 51.37 74.31 62.50 39.66 76.50 51.73 56.36 + ORPO 54.18 74.98 63.46 54.83 77.06 51.18 60.14 + SimPO 53.92 74.48 62.76 38.07 76.56 51.93 56.57 + IFT 54.69 75.08 63.20 57.64 77.27 51.78 60.92
Table 5: Evaluation on Open-LLM Leaderboard when fine-tuning with UltraFeedback-60k.
Method Reference Data Alpaca-Eval Alpaca-Eval-2 pairwise volume win-rate lc win-rate win-rate lc win-rate Gemma-2B      fine-tuning with UltraFeedback-60k + SFT 120k 36.53 30.28 0.99 0.57 + DPO 120k 3.13 1.18 0.13 0.23 + TDPO 120k 2.14 0.70 0.25 0.10 + ORPO 120k 36.62 34.23 1.12 0.59 + SimPO 120k 4.48 2.42 0.13 0.15 + IFT 120k 36.74 39.33 1.61 1.23      fine-tuning with Gemma2-UltraFeedback-armnorm-60k + SFT 120k 39.33 32.36 0.86 0.69 + DPO 120k 2.83 0.81 0.00 0.00 + TDPO 120k 2.41 0.60 0.00 0.00 + ORPO 120k 43.46 34.19 2.06 1.21 + SimPO 120k 3.24 1.07 0.00 0.00 + IFT 120k 51.23 37.76 2.14 1.33
Table 6: Evaluation on LLM-based Benchmarks when fine-tuning with UltraFeedback-60k.
Method ARC ARC-Gen MMLU TruthfulQA WinoGrande GSM8K Avg. Mistral-7B 53.07 73.04 59.14 45.29 77.58 38.89 54.79 + SFT 56.49 74.00 60.44 55.57 77.90 42.84 58.65 + IFT 56.74 74.15 60.49 57.65 78.45 44.73 59.61 + IFT with noisy lambda 61.60 76.53 61.11 57.03 77.43 45.64 60.56
Table 7: Evaluation on Open-LLM Leaderboard when fine-tuning with UltraFeedback-60k.